MPTO(1) MPTO(1) NAME mpto−extractlabelsinMetaPostpicturesfortypesetting SYNOPSIS mpto[-tex|-troff]mpfile DESCRIPTION Thismanualpageisnotmeanttobeexhaustive.Thecompletedocumentationforthisversionof TEXcanbefoundintheinfofileormanualWeb2C:ATeXimplementation. Stripbtex..etexandverbatimtex...etexpartsfromMetaPostinputfilempfile,convertingtoeither TEXorTroff(TEXbydefault).Thegeneratedcodeisdumpedonstandardoutput. mptoistypicallycalledbymakempx,notfromthecommandline. OPTIONS mptoacceptsthefollowingoptions: --helpPrinthelpmessageandexit. -texAssumeinputfilecontainsTEXcommands. -troffAssumeinputfilecontainsTroffcommands. --version Printversioninformationandexit. SEEALSO mpost(1),makempx(1). Web2C7.32x 4January1998 1